OpenStudy (anonymous):

Find all values of k for which the trinomial can be factored. a) x^2 +kx - 15 b) 3x^2 - x + k

OpenStudy (freckles):

what times what is -15?

OpenStudy (freckles):

there are multiple answers to what I'm asking

OpenStudy (jango_in_dtown):

first one -15= 3*-5 and -3*5 now k is the sum of 3 and -5 or sum of -3 and 5 so k is -2 or 2

OpenStudy (jango_in_dtown):

second one the product is 3k and the sum is -1 so 3+k=-1 implies k=-4

OpenStudy (freckles):

-15 can also be written as -1*15 or 1*(-15) @jango_IN_DTOWN

OpenStudy (jango_in_dtown):

@freckles oops, yeah then add them.;)

OpenStudy (netlopes1):

Yes, jango_IN_DTOWN, pefect..... K = 2 or k + -2

OpenStudy (anonymous):

Thanks guys so for B) 3x^2 - x + k k = 2 or k = -2

OpenStudy (anonymous):

and for A) 1 and -15 ?

OpenStudy (anonymous):

could you help explain please

OpenStudy (freckles):

for number 1) your coefficient of x^2 is just 1... so all you need to do is look at what two numbers multiply to be -15 and then just add those numbers to find k. you should get 4 answers here. for number 2) your coefficient of x^2 is something other than 1... so you need to find what two factors of a*c have product a*c and sum b. for number 2 let's say the product of 3k is m*n such that m+n=b where b=-1 in this case so we have 3k=mn and m+n=-1 solve m+n=-1 for either m or n. I choose to solve for m giving me m=-n-1 or m=-(n+1) plugging this into 3k=mn we get 3k=-n(n+1) This means 3k needs to be chosen to be negative product of 2 consecutive integers so 3k=-n(n+1) means 3 could be -n and k could be n+1 or 3k=-n(n+1) means 3 could be n and k could be -(n+1) or 3k=-n(n+1) means 3 could be n+1 and k could be -n or 3k=-n(n+1) means 3 could be -(n+1) and k could be n so of the 4 cases here...playing with these you will only see there actually only 2 possible answers for B

OpenStudy (freckles):

so summary you should have 4 answers for A and 2 answers for B